Show the series Σ x/(1+n^2.x^2 is uniformly convergenr in [ᾰ,1], for any ᾰ>1.


1
Expert's answer
2021-06-29T14:42:18-0400

ANSWER.

Since for each "n\in N" and "x\in \left[ 1,\alpha \right]" , we have "1+{ n }^{ 2 }{ x }^{ 2 }\ge { n }^{ 2 }\ ,\ x\le \alpha" ,then

"0\le \frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } \le \frac { \alpha }{ { n }^{ 2 } }" .

Setting "{ f }_{ n }(x)=\frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } ,\ { M }_{ n }=\frac { \alpha }{ { n }^{ 2 } }" we have for each "n\in N" and "x\in \left[ 1,\alpha \right]" : "\left| { f }_{ n }(x) \right| ={ f }_{ n }(x)\le { M }_{ n }" . The series "\sum _{ n=1 }^{ \infty }{ { M }_{ n }\ } =\alpha \sum _{ n=1 }^{ \infty }{ \frac { 1 }{ { n }^{ 2 } } }" converges, because the series "\sum _{ n=1 }^{ \infty }{ \frac { 1 }{ { n }^{ 2 } } }" is p-series (p=2). By the Weierstrass M-Test , the series

"\sum _{ n=1 }^{ \infty }{ { f }_{ n }(x)= } \sum _{ n=1 }^{ \infty }{ \frac { x }{ 1+{ n }^{ 2 }{ x }^{ 2 } } }"

converges uniformly in "\left[ 1,\alpha \right]" .